# Don’t Mesh with Me: Generating Constructive Solid Geometry Instead of Meshes by Fine-Tuning a Code-Generation LLM

### v1

@@ Abstract
While recent advancements in machine learning, such as LLMs, are revolutionizing software development and creative industries, they have had minimal impact on engineers designing mechanical parts, which remains largely a manual process. Existing approaches to generate 3D geometry most commonly use meshes as a 3D representation. While meshes are suitable for assets in video games or animations, they lack sufficient precision and adaptability for mechanical engineering purposes. This paper introduces a novel approach for the generation of 3D geometry that generates surface-based Constructive Solid Geometry (CSG) by leveraging a code-generation LLM. First, we create a dataset of 3D mechanical parts represented as code scripts by converting Boundary Representation geometry (BREP) into CSG-based Python scripts. Second, we create annotations in natural language using GPT-4. The resulting dataset is used to fine-tune a code-generation LLM. The fine-tuned LLM can complete geometries based on positional input and natural language in a plausible way, demonstrating geometric understanding.
@@ 1 Introduction
In January 2023, ChatGPT [28] reached 100 million monthly active users just two months after launch, setting a record for the fastest-growing user base [15]. The 2024 Stack Overflow Developer Survey found that 62% of developers use AI tools in their workflow, with 82% using them for writing code [40]. Tools like ChatGPT [28] and GitHub Copilot [10] help developers generate or expand code based on text input, speeding up development and extending coding knowledge. Mechanical engineers use CAD software to design 3D parts, but the process remains manual. An AI-driven tool for CAD could accelerate development and allow even non-experts, such as product owners or managers, to create parts through natural language input.
@@
Mechanical parts are mostly made of a combination of basic primitives, such as cylinders, cubes or spheres. Directly representing e.g. a cylinder by a limited set of parameters such as height, radius and position is way more efficient than approximating it using a triangle mesh. It also gives the model precision and makes modifications easier. This type of 3D representation is called boundary representation (BREP) and is commonly used for CAD software. DeepCAD [47] is an approach that can generate CAD sequences, however it is able to generate only random mechanical parts and it has no way of controlling the generation via inputs like natural language. So far, Text2CAD [18] is the only approach that is able to generate CAD sequences from text descriptions. This is achieved by training a transformer architecture on a specific dataset of sketch-based CAD sequences created by Onshape [27] users. This means that the model is limited by data from specific CAD software.
@@
To the best of our knowledge, the resulting model is the first deep learning model specifically trained to generate CSG geometry, and the first fine-tuned LLM that can complete 3D geometry following text instructions.
@@ 2 Related Work
Point cloud generation has seen significant advances through deep generative approaches. Early work leveraging GANs demonstrated promising results in capturing complex 3D geometric structures [37, 43]. Flow-based methods like PointFlow [48] enabled precise generation control through invertible transformations, while gradient field learning [2] offered new shape generation approaches. Autoencoders advanced the field through FoldingNet’s grid deformation [49] and adversarial architectures [51], delivering compact yet accurate point cloud representations. Diffusion-based models have emerged as a powerful alternative, offering improved quality and control over the generation process [52, 26, 22].
@@
Some shape generation methods use Voxel Grids to represent 3D geometry [1, 5, 46, 41]. A notable advancement in voxel grid generation is the application of autoencoders for model variation within a latent space, as presented by Schönhof et al. [35]. By encoding voxel representations into a compact latent space, autoencoders facilitate the generation of diverse voxelized forms through latent space manipulation. Further extending the generative capabilities, 3D-GANs [45] (Generative Adversarial Networks) have been employed to produce voxel grids from a learned distribution of shapes. The integration of natural language with voxel grid generation has been realized through CLIP-Forge [34].
@@
The field of generative models has made progress in the creation of mesh-based geometry by manipulating latent spaces [11, 4, 25, 13]. Autoencoders have been utilized to enable model variation, as explored by Umetani [42]. The area of text-to-mesh generation has made significant progress due to contributions such as the work of Liu et al. [21], which use textual descriptions to generate three-dimensional meshes. Techniques like ShapeCrafter [8] and MeshGPT [38] further extend this capability, allowing direct mesh manipulation using text. MeshGPT specifically allows the generation of 3D geometry with sharp features.
@@
Neural representations of 3D geometry, referred to as neural implicit fields or neural fields encode the geometry of an object or scene as a continuous function represented by a neural network [24, 30]. Building on this foundation, several works have explored using neural fields for generative tasks, such as generating new 3D shapes [9, 31] or novel views of existing scenes [39]. These generative neural field approaches offer the potential for flexible and controllable 3D content creation, with applications in areas like computer graphics, virtual reality, and design.
@@
Approaches such as DeepCAD [47] and Simcurl [6] use a transformer based architecture to generate CAD sequences. DeepCAD, in particular, is able to perform two primary functions: autoencoding existing CAD designs and generating random shapes. Even though these capabilities have limited use cases, they display the model’s ability to generate plausible geometry.
@@ 3 Method
The resulting Python code pair, as shown in Fig. 3, consists of the first part of a 3D model as LLM input (Fig. 3(a)) and the second part as LLM target output (Fig. 3(b)). In addition to the 3D model, the input contains all surface variables that are reused in the output as well as a textual description of the output.
@@ 4.4 Text annotations
Tab. 2 shows the accuracy of the text annotations generated. For every number of cells, 100 text annotations were evaluated by hand and categorized into one of the following categories:

### v2

@@ Abstract
While recent advancements in machine learning, such as LLMs, are revolutionizing software development and creative industries, they have had minimal impact on engineers designing mechanical parts, which remains largely a manual process. Existing approaches to generating 3D geometry most commonly use meshes as a 3D representation. While meshes are suitable for assets in video games or animations, they lack sufficient precision and adaptability for mechanical engineering purposes. This paper introduces a novel approach for the generation of 3D geometry that generates surface-based Constructive Solid Geometry (CSG) by leveraging a code-generation LLM. First, we create a dataset of 3D mechanical parts represented as code scripts by converting Boundary Representation geometry (BREP) into CSG-based Python scripts. Second, we create annotations in natural language using GPT-4. The resulting dataset is used to fine-tune a code-generation LLM. The fine-tuned LLM can complete geometries based on positional input and natural language in a plausible way, demonstrating geometric understanding.
@@ 1 Introduction
In January 2023, ChatGPT [27] reached 100 million monthly active users just two months after its launch, setting a record for the fastest-growing user base [14]. The 2024 Stack Overflow Developer Survey found that 62% of developers use AI tools in their workflow, with 82% using them for writing code [39]. Tools like ChatGPT [27] and GitHub Copilot [9] help developers generate or expand code based on text input, speeding up development and extending coding knowledge. Mechanical engineers use CAD software to design 3D parts, but the process remains manual. An AI-driven tool for CAD could accelerate development and allow even non-experts, such as product owners or managers, to create parts through natural language input.
@@
Mechanical parts are mostly made of a combination of basic primitives, such as cylinders, cubes or spheres. Directly representing e.g. a cylinder by a limited set of parameters such as height, radius and position is way more efficient than approximating it using a triangle mesh. It also gives the model precision and makes modifications easier. This type of 3D representation is called boundary representation (BREP) and is commonly used for CAD software. DeepCAD [46] is an approach that can generate CAD sequences, however it is able to generate only random mechanical parts and it has no way of controlling the generation via inputs like natural language. So far, Text2CAD [17] is the only approach that can generate CAD sequences from text descriptions. This is achieved by training a transformer architecture on a specific dataset of sketch-based CAD sequences created by Onshape [26] users. This means that the model is limited by data from specific CAD software.
@@
To the best of our knowledge, the resulting model is the first deep learning model specifically trained to generate CSG geometry, and the first LLM that is fine-tuned to complete 3D geometry following text instructions.
@@ 2 Related Work
Point cloud generation has progressed through GANs [36, 42], invertible flows [50], gradient fields [2], autoencoders [51, 53], and more recently, diffusion models [54, 25, 21], each improving control or representation fidelity.
@@
Some shape generation methods use Voxel Grids to represent 3D geometry [1, 5, 45, 40]. A notable advancement in voxel grid generation is the application of autoencoders for model variation within a latent space, as presented by Schönhof et al. [34]. By encoding voxel representations into a compact latent space, autoencoders facilitate the generation of diverse voxelized forms through latent space manipulation. Further extending the generative capabilities, 3D-GANs [44] (Generative Adversarial Networks) have been employed to produce voxel grids from a learned distribution of shapes.
@@
The field of generative models has made progress in the creation of mesh-based geometry by manipulating latent spaces [10, 4, 24, 12]. Autoencoders have been utilized to enable model variation, as explored by Umetani [41]. The field of text-to-mesh generation has seen notable advancements, particularly through approaches that leverage implicit field representations for mesh extraction. Prominent examples include the works of Liu et al. [20] and ShapeCrafter [7], both of which utilize natural language input to guide 3D shape generation. In contrast, MeshGPT [37] employs a transformer-based architecture to generate 3D meshes directly, bypassing the need for implicit field modeling. Notably, MeshGPT is capable of producing geometry with sharp and well-defined features.
@@
Neural representations of 3D geometry, referred to as neural implicit fields or neural fields encode the geometry of an object or scene as a continuous function represented by a neural network [23, 29]. Building on this foundation, several works have explored using neural fields for generative tasks, such as generating new 3D shapes [8, 30, 33] or novel views of existing scenes [38]. These generative neural field approaches offer the potential for flexible and controllable 3D content creation, with applications in areas like computer graphics, virtual reality, and design.
@@
BREP Generation
BrepGen [49] is a recent diffusion-based method that directly generates CAD models as boundary representations (BREPs). It introduces a hierarchical tree structure where each node encodes geometric primitives, and topology is captured implicitly through node duplication. BrepGen generates 3D geometry suitable for CAD applications. However, it is not capable of generating geometry based on natural language.
@@
Approaches such as DeepCAD [46], SkexGen [47] and HNC-CAD [48] use transformer based architectures to generate CAD sequences. DeepCAD primarily performs two functions: autoencoding CAD designs and generating random shapes. While their use cases are limited, they showcase the model’s ability to produce plausible geometry.
@@ 3 Method
The resulting Python code pair, as shown in Fig. 3, consists of the first part of a 3D model as LLM input (Fig. 3(a)) and the second part as LLM target output (Fig. 3(b)). The geometric input and output contain at least one cell each. In addition to the 3D model, the input contains all surface variables that are reused in the output as well as a textual description of the output.
@@ 4.4 Text annotations
Tab. 2 shows the accuracy of the text annotations generated. For every number of cells, 100 text annotations were human evaluated and categorized into one of the following categories: